OpenStudy (anonymous):

Which statements are true according to the order of operations? Choose exactly two answers that are correct. A. 2 + 5 • 7 = 2 + 35 = 37 B. 10 + (6 ÷ 3) = 10 + 2 = 12 C. 16 – (4 + 7) = 12 + 7 = 19 D. 7 • 6 + 4 = 7 • 10 = 70

OpenStudy (tkhunny):

A. 2 + 5 • 7 = 2 + 35 = 37 What do you think? Did multiplication happen before addition? B. 10 + (6 ÷ 3) = 10 + 2 = 12 What do you think? Did parentheses happen before addition? C. 16 – (4 + 7) = 12 + 7 = 19 What do you think? Did parentheses happen before subtraction? D. 7 • 6 + 4 = 7 • 10 = 70 What do you think? Did multiplication happen before addition?